OpenStudy (mouth4war):

i really really need help with understanding things. could somebody help me?

OpenStudy (ebonygotthejuice):

i could try

OpenStudy (mouth4war):

how do I do this? i need to understand it but its difficult and i dont know how to get the answer or do the work to find the answer.

OpenStudy (mouth4war):

@ebonygotthejuice

OpenStudy (ebonygotthejuice):

hold on

OpenStudy (ebonygotthejuice):

@dmezzullo

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (mouth4war):

i dont understand that attachment??

OpenStudy (stacey):



OpenStudy (stacey):

Theta is the 49 degree angle, so which side is the hypotenuse and which is the opposite side?

OpenStudy (vespasious):

Okay, Angle C is 49 degrees. Triangles have 180 degrees, so we know that (180-90-49) = 41. That's angle A. To find the height of the wall, we can use tangent, since we're finding an opposite side and we have an adjacent side. To set this up, tan(41) = x/3 3*tan(41) = x x= 2.607. Now that we have the distance from the plank to the wall, we can use Pythagorean formula to find the hypotenuse. c^2=2.607^2+3^2 c^2 = (6.796) + 9 c = sqrt (15.796) = 3.97. Now you have all your side lengths and angles. To recap, <A = 41 <B = 90 <C = 49 (given) A->B = 3ft (given) B->C = 2.607 C->A = 3.97 Good luck!